# Couples can be tractable: New algorithms and hardness results for the Hospitals / Residents problem with Couples

### v2

@@ Abstract
Our polynomial-time solvability results greatly expand the class of known tractable instances of hrc and provide additional evidence as to why long-standing entry-level labour markets that allow couples such as the National Resident Matching Program remain successful to this day.
@@ 1.1 Background
The Stable Marriage problem, and its many-to-one extension, the Hospitals / Residents problem (hr), are well-studied and central problems in Computational Social Choice, Computer Science, Game Theory and Economics, with numerous applications including in entry-level labour markets, school choice and higher education allocation [22]. In the medical sphere, centralised matching schemes that assign aspiring junior doctors to hospitals operate in many countries. One of the largest and best known examples is the National Resident Matching Program (NRMP) in the US [36], which had just under 43,000 applicants for just over 40,000 residency positions in 2023 [37]. There are analogous matching schemes for junior doctor allocation in Canada [38], France [39] and Japan [40].
@@
hr is also known to have very nice mathematical structure. In particular, the famous Rural Hospitals Theorem states that, for an arbitrary hr instance $I$, any doctor that is assigned in one stable matching in $I$ is assigned in all stable matchings in $I$, moreover any hospital that is under-subscribed in some stable matching in $I$ is assigned exactly the same set of doctors in every stable matching in $I$ [30, 13, 31].
@@ 1.2 Related work
Roth [30] considered stability in the hrc context although did not define the concept explicitly. Whilst Gusfield and Irving [14] defined stability formally in hrc, their definition neglected to deal with the case that both members of a couple may wish to be assigned to the same hospital simultaneously. There are discussions of this point in [5] and [22, Section 5.3.2]. McDermid and Manlove [25] provided a stability definition for hrc that does deal with this possibility. Other such definitions have been proposed in the literature [4, 21], and a detailed comparative study of the effects of different stability definitions for hrc was carried out by Delorme et al. [9].
@@
Given the prevalence of NP-hardness results for hrc, and the scarcity of polynomial-time algorithms, heuristics have been applied to the problem (see [22, Section 5.3,3] for a survey) as well as approaches based on parameterised complexity and local search [24, 4], integer programming [6, 9], constraint programming [23] and SAT solving [10].
@@ 1.3 Our contributions
Our main result is a novel and surprising polynomial-time algorithm to find a near-feasible stable matching in an hrc instance if the couples’ preferences are sub-responsive and sub-complete. In this paper, our notion of near-feasibility is based on modifying the capacities of each hospital by at most 1. This strengthens Nguyen and Vohra’s result [28], albeit for a special case of hrc, in two ways: (i) capacities are varied by at most 1 rather than 2, and (ii) we provide a polynomial-time algorithm to produce the desired outcome. This result is established via a reduction to the Stable Fixtures problem (sf), a non-bipartite many-to-many generalisation of hr [18]. Our algorithm has other nice properties: for example, it can indeed be very beneficial for a couple to apply together, as it can happen that by applying alone, only one of them would have been accepted, but by applying together, both of them does. We illustrate this property with an example.
@@ 2.1 Notation, terminology and problem models
We now define the concept of a blocking pair in hrc, adopting the definition from [25].
@@
A couple $C_{i}=(c_{2i-1},c_{2i})\in C$ forms a blocking pair of $M$ with a hospital pair $(h_{j},h_{k})\in{(H\cup\{\emptyset\})\times(H\cup\{\emptyset\})}$ if $(h_{j},h_{k})\in A(C_{i})$, $(h_{j},h_{k})\succ_{C_{i}}M(C_{i})$, and either:
@@
In hrc, the problem is to find a stable matching or report that none exists. An important special case of hrc arises when the set $H^{\prime}$ of acceptable hospitals for the couple members, i.e., $H^{\prime}=\bigcup_{c_{i}\in C^{\prime}}A(c_{i})$, can be partitioned into two sets $H_{1},H_{2}$, such that for each couple $C_{i}\in C$, each acceptable pair $(h_{j},h_{k})\in A(C_{i})$ satisfies $h_{j}\in H_{1}$ and $h_{k}\in H_{2}$. We refer to this restriction of hrc as hrc-dual market (corresponding to a Dual Market).
@@
  • separable, if $A(C_{i})=(A(c_{2i-1})\cup\{\emptyset\})\times(A(c_{2i})\cup\{\emptyset\})\setminus\{(\emptyset,\emptyset)\}$ (that is, it is allowed to match only one member of the couple in a feasible matching),
  • half-separable if $A(C_{i})=A(c_{2i-1})\times(A(c_{2i})\cup\{\emptyset\})$ or $(A(c_{2i-1})\cup\{\emptyset\})\times A(c_{2i})$ (that is, it is allowed to leave at most one member of the couple unmatched, but this can only ever be one designated member),
@@
  • of type-b, if it is connected, $A(c_{2i-1})\cap A(c_{2i})=\{h\}$ and if $c_{2i-1}\succ_{h}c_{2i}$, then $A(c_{2i})=\{h\}$, otherwise $A(c_{2i-1})=\{h\}$ (that is, the worse member for $h$ only applies to $h$)
  • of type-c, if it is connected, $A(c_{2i-1})\cap A(c_{2i})=\{h\}$ and for any $h\neq h^{\prime}\in A(c_{2i-1})$, $h^{\prime}\succ_{c_{2i-1}}h$ and for any $h\neq h^{\prime\prime}\in A(c_{2i})$, $h^{\prime\prime}\succ_{c_{2i}}h$, (that is, $h$ is the worst hospital for both members of the couple).
@@
We will show that hrc is solvable in polynomial time for couples of type-a, -b or -c. Moreover for Dual Markets, every couple is necessarily of type-a. However, every couple being of type-a is more general than hrc-dual market, as it only means that for each couple $C_{i}$, the acceptable hospitals for each member of $C_{i}$ are two disjoint sets. To illustrate this with a simple example, imagine that there are three hospitals $h_{1},h_{2},h_{3}$ and three couples $C_{1},C_{2},C_{3}$. $C_{1}$ applies only to $(h_{1},h_{2})$, $C_{2}$ applies only to $(h_{2},h_{3})$, and $C_{3}$ applies only to $(h_{3},h_{1})$. It is easy to see that every couple is of type-a. However, the market is not dual, as for any possible partition of $\{h_{1},h_{2},h_{3}\}$ into two parts, there will be an application where both hospitals are from the same part.
Our results also imply a nice corollary about a stable matching problem in multigraphs, which we now define. The setting is a multigraph $G=(V,E)$ where loops are also allowed. Each node $v\in V$ has a strict preference list $\succ_{v}$ over the edges incident to $v$. We are also given a capacity $b(v)$ for each $v\in V$. A b-matching is a set $M\subseteq E$ such that $|M(v)|\leq b(v)$ for each $v\in V$, where $M(v)$ is the set of edges in $M$ that are incident to $v$. In this multigraph b-matching model, stability is defined as follows.
@@
The Stable Fixtures problem (sf) [18], also called the Stable b-Matching problem, is defined similarly, but the underlying graph $G$ is a simple graph, so it is not allowed to have loops or parallel edges, and hence the notion of stability also simplifies. The generalisation of sf  where parallel edges are allowed, but loops are not, has been studied by Cechlárová and Fleiner [7], who called this problem Stable Multiple Activities. We note that our new problem variant does not reduce straightforwardly to Stable Multiple Activities. The main difficulty lies in the fact that for a loop edge to block, it needs two worse edges or free positions at a node, so if a node is filled with better edges than the loop, except for one worse edge, then no endpoint of the loop is dominated, but we still do not allow it to block.
@@ 2.2 Motivation for couple types
We also note the important special case of hrc in which each couple applies only to one pair of hospitals. Then as per the proof of Corollary 14, each couple is type-a or -b. On the other hand, hrc is NP-hard even if each couple finds acceptable only two hospital pairs [6]. This helps to push forward our understanding of the frontier between polynomial-time solvability and NP-hardness for restrictions of hrc. Indeed, couple structures that do not make the problem NP-hard may need to be very specific, as Theorem 18 shows that hrc remains NP-hard even with sub-responsive and sub-complete couples, such that each member of a couple applies to only two hospitals.
As a more theoretical application, types-a, -b and -c also arise when we model a new multigraph extension of sf where loop edges are allowed. Due to space constraints, this problem is not discussed in the current version of this paper. Finally, we note that even in cases where there may be different types of couples (beyond those defined in Definitions 2 and 3), it may speed up heuristics and IP techniques to first compute (in polynomial time) a stable solution restricted to the single doctors and the couples of types-a, -b and -c.
@@ 3 Polynomial-time algorithms
Given an hrc instance where each couple’s preferences are sub-responsive and sub-complete, then there is always a near-feasible capacity vector $\textbf{q}^{\prime}$ where $|\textbf{q}^{\prime}-\textbf{q}|_{\infty}\leq 1$, such that there is a stable matching with respect to the modified capacities. Furthermore, these modified capacities and the stable matching can be found in $\mathcal{O}(m)$ time, where $m$ denotes the total length of the preference lists of the hospitals.
@@
Let $I=(D,C,H,\mathbf{q},\succ)$ be an instance of hrc where the couples’ preferences are sub-responsive and sub-complete. We transform $I$ to a sf instance $I^{\prime}$, and start by describing the nodes and the capacities in $I^{\prime}$.
@@
  • Let $c_{k}$ be a member of a non-half-separable couple. The preferences of $c_{k}$ in $I^{\prime}$ are then created such that he/she ranks $a_{k}$ first (if it exists), followed by the hospitals in $A(c_{k})$ according to $\succ_{c_{k}}$, and $b_{k-1}$ or $b_{k+1}$ last (depending on the parity of $k$, again, if it exists).
@@
In $M_{f}$, each node is incident to either 0 or two incident fractional edges.
@@
If $c_{k}$ is a member of a half-separable couple who cannot be assigned alone, then we show that $c_{k}$ is assigned only if his partner $c_{k^{\prime}}$ is too. Indeed, if $c_{k^{\prime}}$ is not matched to a hospital, but $c_{k}$ is, then $c_{k^{\prime}}$ is not matched with weight 1 in $M_{f}$, and as the edge $(c_{k},c_{k^{\prime}})$ is the best for $c_{k}$, we get that it blocks $M_{f}$, contradiction.
@@
Suppose that a couple $C_{i}\in C$ blocks $M$ in $I$ with a pair of hospitals $(h_{j},h_{k})\in A(C_{i})$. The following argument holds regardless of whether $h_{j}=h_{k}$ or not. By the assumption that the couple’s preferences are sub-responsive, it follows that either $C_{i}$ is unmatched in $M_{f}$ (in which case one of the members cannot be matched with edges better than his hospital partners with weight 1 in $M_{f}$), or one of its members $c_{2i-1}$, $c_{2i}$ is at a worse hospital according to his ranking. Suppose by symmetry that this member is $c_{2i}$. By the construction of $M$, it follows that $M_{f}(c_{2i},h_{k})=0$ must hold. But the edge $(c_{2i},h_{k})$ did not block $M_{f}$ in $I^{\prime}$, so $h_{k}$ was saturated in $M_{f}$ with better partners than $c_{2i}$, all of them being doctors. Hence, we obtain again by our construction that $h_{k}$ is full in $M$ with better doctors than $c_{2i}$, so $C_{i}$ cannot block $M$ in $I$ with $(h_{j},h_{k})$ after all.
Suppose a couple $C_{i}$ blocks with one hospital and an application $(h_{i},\emptyset)$ or $(\emptyset,h_{j})$. In this case, $C_{i}$ must be separable or half-separable. Suppose by symmetry that $c_{2i}$ is the one who gets matched to a hospital in the blocking coalition. Then, if $C_{i}$ is half-separable, then $c_{2i}$ must be the member who can be assigned alone, so he ranks $c_{2i-1}$ last. Hence, in any case, $c_{2i}$ cannot be matched to better partners in $M_{f}$ than $h_{j}$, and $M_{f}(c_{2i},h_{j})=0$. But $(c_{2i},h_{j})$ does not block $M_{f}$, so $h_{j}$ was saturated with better doctors in $M_{f}$, so it still is in $M$, contradiction.
@@
By the result of Nguyen and Vohra [28], we know that there is always a near-feasible stable matching in any hrc instance, if we can change the capacities by 2 instead of 1. However, their approach to finding such a near-feasible stable matching starts by constructing a fractional stable solution, which by a recent paper of Csáji [8] is shown to be computationally hard (PPAD-hard). Therefore, even though the existence is guaranteed, there is no efficient algorithm yet that can find a near-feasible stable solution in an arbitrary hrc instance, and even if such an algorithm exists, it cannot start by finding a stable fractional solution as per the approach of Nguyen and Vohra.
In the hrc problem, especially with sub-responsive and sub-complete preferences, it might be unclear as to why it is beneficial for a couple to apply together to pairs of hospitals. Indeed, any stable matching would be stable even if they would have applied separately, however, by applying together it is possible that neither of them gets matched in the end, whereas one of them could have been matched if they had applied separately. However, a very nice feature of our algorithm as given in the proof of Theorem 5 is that now it may be very beneficial for a couple to apply together. We demonstrate this via a simple example.
@@
We prove our statement by creating a sf instance using a construction that extends the polynomial-time reduction given in the proof of Theorem 5. Reindex the couples $C_{i}$ of type-b and c, such that $c_{2i-1}$ is the worse member of $C_{i}$ for the common hospital $h$. We will use this fact to make the analysis simpler. We proceed by describing our reduction. Let $I=(D,C,H,\mathbf{q},\succ)$ be an instance of hrc satisfying the conditions of the theorem.
@@
    • If $c_{k}$ is a member of a couple, then each occurrence of $h_{j}\in A(c_{k})$ is replaced by $h_{j}^{1}\succ_{c_{k}}^{\prime}h_{j}^{2}$ in $\succ_{c_{k}}^{\prime}$.
@@
First note that by adding dummy single doctors that find each hospital acceptable, and each hospital considers them worst, we can assume that each hospital is filled in $M$. (For the algorithm, we do not need to add these dummy doctors, it just makes the analysis simpler.) We create a matching $M^{\prime}$ in $I^{\prime}$ as follows.
For each hospital $h_{j}$, let $M^{1}(h_{j})$ be the best $q_{j}-1$ doctors at $h_{j}$ and $M^{2}(h_{j})$ be the worst doctor at $h_{j}$. Then in $M^{\prime}$, we match the nodes corresponding to the doctors in $M^{1}(h_{j})$ to $h_{j}^{1}$ and the doctor in $M^{2}(h_{j})$ to $h_{j}^{2}$. Note that it cannot happen that we match a doctor $c_{k}$, who is a member of a couple, to $h_{j}^{2}$ if the edge $(c_{k},h_{j}^{2})$ does not exist: indeed, in that case $c_{k}$ cannot be matched alone and his partner $c_{k^{\prime}}$ only applies to $h_{j}$ and is worse for $h_{j}$, so $c_{k}$ cannot be the worst doctor in $h_{j}$. Note that as each hospital $h_{j}$ was filled in $M$, every node $h_{j}^{1},h_{j}^{2}$ is filled by $M^{\prime}$.

### v3

@@ Abstract
Our polynomial-time solvability results greatly expand the class of known tractable instances of hrc and provide a useful tool for designing better and more efficient mechanisms in the future.
@@ 1.1 Background
The Stable Marriage problem, and its many-to-one extension, the Hospitals / Residents problem (hr), are well-studied and central problems in Computational Social Choice, Computer Science, Game Theory and Economics, with numerous applications including in entry-level labour markets, school choice and higher education allocation [23]. In the medical sphere, centralised matching schemes that assign aspiring junior doctors to hospitals operate in many countries. One of the largest and best known examples is the National Resident Matching Program (NRMP) in the US [37], which had just under 45,000 applicants for just over 41,500 positions in the Main Residency Match in 2024 [38]. There are analogous matching schemes for junior doctor allocation in Canada [39], France [40] and Japan [41].
@@
hr is also known to have very nice mathematical structure. In particular, the famous Rural Hospitals Theorem states that, for an arbitrary hr instance $I$, any doctor that is assigned in one stable matching in $I$ is assigned in all stable matchings in $I$, moreover any hospital that is undersubscribed in some stable matching in $I$ is assigned exactly the same set of doctors in every stable matching in $I$ [31, 14, 32].
@@ 1.2 Related work
Roth [31] considered stability in the hrc context although did not define the concept explicitly. Whilst Gusfield and Irving [15] defined stability formally in hrc, their definition neglected to deal with the case that both members of a couple may wish to be assigned to the same hospital simultaneously. There are discussions of this point in [5] and [23, Section 5.3.2]. McDermid and Manlove [26] provided a stability definition for hrc that does deal with this possibility. Other such definitions have been proposed in the literature [4, 22], and a detailed comparative study of the effects of different stability definitions for hrc was carried out by Delorme et al. [10].
@@
Given the prevalence of NP-hardness results for hrc, and the scarcity of polynomial-time algorithms, heuristics have been applied to the problem (see [23, Section 5.3,3] for a survey) as well as approaches based on parameterised complexity and local search [25, 4], integer programming [6, 10], constraint programming [24] and SAT solving [11].
@@ 1.3 Our contributions
Our main result is a novel and surprising polynomial-time algorithm to find a near-feasible stable matching in an hrc instance if the couples’ preferences are sub-responsive and sub-complete. In this paper, our notion of near-feasibility is based on modifying the capacities of each hospital by at most 1. This strengthens Nguyen and Vohra’s result [29], albeit for a special case of hrc, in two ways: (i) capacities are varied by at most 1 rather than 2, and (ii) we provide a polynomial-time algorithm to produce the desired outcome. This result is established via a reduction to the Stable Fixtures problem (sf), a non-bipartite many-to-many generalisation of hr [19]. Our algorithm has other nice properties: for example, it can indeed be very beneficial for a couple to apply together, as it can happen that by applying alone, only one of them would have been accepted, but by applying together, both of them are. We illustrate this property with an example.
@@ 2.1 Notation, terminology and problem models
We now define the concept of a blocking pair in hrc, adopting the definition from [26]. While the definition seems complicated, the underlying idea of it is just to consider the possible cases for single doctors, couples and hospitals to form a mutually beneficial deviation from the matching.
@@
A couple $C_{i}=(c_{2i-1},c_{2i})\in C$ forms a blocking pair of $M$ with a hospital pair $(h_{j},h_{k})\in{(H\cup\{\emptyset\})\times(H\cup\{\emptyset\})}$ if $(h_{j},h_{k})\in A(C_{i})$, $C_{i}$ is unmatched or $(h_{j},h_{k})\succ_{C_{i}}M(C_{i})$, and either:
@@
In hrc, the problem is to find a stable matching or report that none exists. An important special case of hrc arises when the set $H$ of hospitals can be partitioned into two sets $H_{1},H_{2}$, such that for each couple $C_{i}\in C$, each acceptable pair $(h_{j},h_{k})\in A(C_{i})$ satisfies $h_{j}\in H_{1}$ and $h_{k}\in H_{2}$, and for each single doctor $d_{i}$, either $A(d_{i})\subseteq H_{1}$ or $A(d_{i})\subseteq H_{2}$. We refer to this restriction of hrc as hrc-dual market (corresponding to a Dual Market).
@@
  • separable, if $A(C_{i})=((A(c_{2i-1})\cup\{\emptyset\})\times(A(c_{2i})\cup\{\emptyset\}))\setminus\{(\emptyset,\emptyset)\}$ (that is, it is allowed to match only one member of the couple in a feasible matching),
  • half-separable if $A(C_{i})=A(c_{2i-1})\times(A(c_{2i})\cup\{\emptyset\})$ or $A(C_{i})=(A(c_{2i-1})\cup\{\emptyset\})\times A(c_{2i})$ (that is, it is allowed to leave at most one member of the couple unmatched, but this can only ever be one designated member),
@@
  • of type-b, if it is connected, $A(c_{2i-1})\cap A(c_{2i})=\{h\}$ for some $h\in H$, and if $c_{2i-1}\succ_{h}c_{2i}$, then $A(c_{2i})=\{h\}$, otherwise $A(c_{2i-1})=\{h\}$ (that is, the worst member of the couple for $h$ only applies to $h$)
  • of type-c, if it is connected, $A(c_{2i-1})\cap A(c_{2i})=\{h\}$ for some $h\in H$, and for any $h^{\prime}\in A(c_{2i-1})\setminus\{h\}$, $h^{\prime}\succ_{c_{2i-1}}h$ and for any $h^{\prime\prime}\in A(c_{2i})\setminus\{h\}$, $h^{\prime\prime}\succ_{c_{2i}}h$, (that is, $h$ is the worst hospital for both members of the couple).
@@
We will show that hrc is solvable in polynomial time for couples of type-a, -b or -c. Moreover for Dual Markets, every sub-responsive and sub-complete couple is necessarily of type-a. However, every couple being of type-a is more general than hrc-dual market, as it only means that for each couple $C_{i}$, the acceptable hospitals for each member of $C_{i}$ are two disjoint sets. To illustrate this with a simple example, imagine that there are three hospitals $h_{1},h_{2},h_{3}$ and three couples $C_{1},C_{2},C_{3}$. $C_{1}$ applies only to $(h_{1},h_{2})$, $C_{2}$ applies only to $(h_{2},h_{3})$, and $C_{3}$ applies only to $(h_{3},h_{1})$. It is easy to see that every couple is of type-a. However, the market is not dual, as for any possible partition of $\{h_{1},h_{2},h_{3}\}$ into two parts, there will be an application where both hospitals are from the same part.
Our results will also imply a nice corollary about a stable matching problem in multigraphs, which we now define. The setting is a multigraph $G=(V,E)$ where loops are also allowed. Each node $v\in V$ has a strict preference list $\succ_{v}$ over the edges incident to $v$. We are also given a capacity $b(v)$ for each $v\in V$. A b-matching is a set $M\subseteq E$ such that $|M(v)|\leq b(v)$ for each $v\in V$, where $M(v)$ is the set of edges in $M$ that are incident to $v$. In this multigraph b-matching model, stability is defined as follows.
@@
The Stable Fixtures problem (sf) [19], also called the Stable b-Matching problem, is defined similarly, but the underlying graph $G$ is a simple graph, so it is not allowed to have loops or parallel edges, and hence the notion of stability also simplifies. The generalisation of sf  where parallel edges are allowed, but loops are not, has been studied by Cechlárová and Fleiner [7], who called this problem Stable Multiple Activities. We note that our new problem variant does not reduce straightforwardly to Stable Multiple Activities. The main difficulty lies in the fact that for a loop edge $e=(u,u)$ to block, it needs two worse edges or free positions at node $u$, so if $u$ is filled with better edges than $e$ except for one worse edge, still it is the case that $e$ is not allowed to block.
@@ 2.2 Motivation for couple types
We also note the important special case of hrc in which each couple applies only to one pair of hospitals. Then as per the proof of Corollary 14 below, each couple is type-a or -b. On the other hand, hrc is NP-hard even if each couple finds acceptable only two hospital pairs [6]. This helps to push forward our understanding of the frontier between polynomial-time solvability and NP-hardness for restrictions of hrc. Indeed, couple structures that do not make the problem NP-hard may need to be very specific, as Theorem 19 below shows that hrc remains NP-hard even with sub-responsive and sub-complete couples, such that each member of a couple applies to only two hospitals.
As a more theoretical application, couples types-a and -b also arise in Stable Multigraph b-Matching, as established in Corollary 16 below. Finally, we note that even in cases where there may be different types of couples (beyond those defined in Definitions 2 and 3), it may speed up heuristics and IP techniques to first compute (in polynomial time) a stable solution restricted to the single doctors and the couples of types-a, -b and -c.
@@ 3 Polynomial-time algorithms
Given an hrc instance where each couple’s preferences are sub-responsive and sub-complete, there is always a near-feasible capacity vector $\textbf{q}^{\prime}$ where $|\textbf{q}^{\prime}-\textbf{q}|_{\infty}\leq 1$, such that there is a stable matching with respect to the modified capacities. Furthermore, these modified capacities and the stable matching can be found in $\mathcal{O}(m)$ time, where $m$ denotes the total length of the preference lists of the hospitals.
@@
Let $I=(D,C,H,\mathbf{q},\succ)$ be an instance of hrc where the couples’ preferences are sub-responsive and sub-complete. We transform $I$ to an sf instance $I^{\prime}$, and start by describing the nodes and the capacities in $I^{\prime}$.
@@
  • Let $c_{k}$ be a member of a connected couple. The preferences of $c_{k}$ in $I^{\prime}$ are then created such that he/she ranks $a_{k}$ first (if it exists), followed by the hospitals in $A(c_{k})$ according to $\succ_{c_{k}}$, and $b_{k-1}$ or $b_{k+1}$ last (depending on the parity of $k$, again, if it exists).
@@
In $M_{f}$, each node is incident to either 0 or two fractional edges.
@@
If $c_{k}$ is a member of a half-separable couple who cannot be assigned alone, then we show that $c_{k}$ is assigned only if his partner $c_{k^{\prime}}$ is too. Indeed, if $c_{k^{\prime}}$ is not matched to a hospital node, but $c_{k}$ is, then $c_{k^{\prime}}$ is not matched with weight 1 in $M_{f}$ to a hospital node, and as the edge $(c_{k},c_{k^{\prime}})$ is the best for $c_{k}$, it follows that it blocks $M_{f}$ in $I^{\prime}$, contradiction.
@@
Suppose that a couple $C_{i}\in C$ blocks $M$ in $I$ with a pair of hospitals $(h_{j},h_{k})\in A(C_{i})$. The following argument holds regardless of whether $h_{j}=h_{k}$ or not. By the assumption that the couple’s preferences are sub-responsive and sub-complete, it follows that either $C_{i}$ is unmatched in $M$, in which case at least one of the members’ nodes is not full in $M_{f}$ with better edges than $(c_{2i-1},h_{j})$ or $(c_{2i},h_{k})$ respectively (in the case of a connected couple one of $\{c_{2i-1},c_{2i}\}$ is not matched to his favourite partner $a_{2i}$ or $a_{2i-1}$ with weight 1 in $M_{f}$, as otherwise $(a_{2i},b_{2i})$ blocks $M_{f}$); or one of its members $c_{2i-1}$, $c_{2i}$ is at a worse hospital in $M$ according to his ranking. In both cases, one of the edges $\{(c_{2i-1},h_{j}),(c_{2i},h_{k})\}$ is not dominated at a couple node in $M_{f}$. Suppose by symmetry it is $(c_{2i},h_{k})$. By the construction of $M$, it follows that $M_{f}(c_{2i},h_{k})=0$ must hold. However, the edge $(c_{2i},h_{k})$ did not block $M_{f}$ in $I^{\prime}$, so $h_{k}$ was saturated in $M_{f}$ with better partners than $c_{2i}$, all of them being doctors. Hence, we obtain again by our construction that $h_{k}$ is full in $M$ with better doctors than $c_{2i}$, so $C_{i}$ cannot block $M$ in $I$ with $(h_{j},h_{k})$.
Suppose a couple $C_{i}$ blocks with one hospital and an application $(h_{i},\emptyset)$ or $(\emptyset,h_{j})$. In this case, $C_{i}$ must be separable or half-separable. Suppose by symmetry that $c_{2i}$ is the one who gets matched to a hospital in the blocking coalition. Then, if $C_{i}$ is half-separable, then $c_{2i}$ must be the member who can be assigned alone, so he ranks $c_{2i-1}$ last. Hence, in any case, $c_{2i}$ cannot be matched to better partners in $M_{f}$ than $h_{j}$, and $M_{f}(c_{2i},h_{j})=0$. However, $(c_{2i},h_{j})$ did not block $M_{f}$, so $h_{j}$ was saturated with better doctors in $M_{f}$, so it still is in $M$, contradiction.
@@
By the result of Nguyen and Vohra [29], we know that there is always a near-feasible stable matching in any hrc instance, if we can change the capacities by 2 instead of 1. However, their approach to finding such a near-feasible stable matching starts by constructing a fractional stable solution, which by a recent paper of Csáji [8] is shown to be computationally hard (PPAD-hard) in general. Therefore, even though the existence is guaranteed, there is no efficient algorithm yet that can find a near-feasible stable solution in an arbitrary hrc instance.
In the hrc problem, especially with sub-responsive and sub-complete preferences, it might be unclear as to why it is beneficial for a couple to apply together to pairs of hospitals. Indeed, any stable matching would be stable even if they would have applied separately. However, a very nice feature of our algorithm as given in the proof of Theorem 5 is that now it may be very beneficial for a couple to apply together, as they can obtain a better match. We demonstrate this via a simple example.
@@
We prove our statement by creating an sf instance using a construction that extends the polynomial-time reduction given in the proof of Theorem 5. Let $I=(D,C,H,\mathbf{q},\succ)$ be an instance of hrc satisfying the conditions of the theorem. Reindex the couples $C_{i}$ of type-b and c, such that $c_{2i-1}$ is the worse member of $C_{i}$ for the common hospital $h$. We will use this fact to make the analysis simpler. We proceed by describing our reduction.
@@
    • First, each occurrence of $h_{j}\in A(c_{k})$ is replaced by $h_{j}^{1}\succ_{c_{k}}^{\prime}h_{j}^{2}$ in $\succ_{c_{k}}^{\prime}$.
@@
First note that by adding dummy single doctors that find each hospital acceptable, and each hospital considers them worst, we can assume that each hospital is filled in $M$. (For the algorithm, we do not need to add these dummy doctors, it just helps to avoid further case distinctions in the analysis, on whether a hospital has worse doctors or free positions). We create a matching $M^{\prime}$ in $I^{\prime}$ as follows.
For each hospital $h_{j}$, let $M^{1}(h_{j})$ be the best $q_{j}-1$ doctors at $h_{j}$ and $M^{2}(h_{j})$ be the worst doctor at $h_{j}$. Then, in $M^{\prime}$, we match the nodes corresponding to the doctors in $M^{1}(h_{j})$ to $h_{j}^{1}$ and the doctor in $M^{2}(h_{j})$ to $h_{j}^{2}$. Note that it cannot happen that we match a doctor $c_{k}$, who is a member of a couple, to $h_{j}^{2}$ if the edge $(c_{k},h_{j}^{2})$ does not exist: indeed, in that case $c_{k}$ cannot be matched alone and his partner $c_{k^{\prime}}$ only applies to $h_{j}$ and is worse for $h_{j}$, so $c_{k}$ cannot be the worst doctor at $h_{j}$. Note that as each hospital $h_{j}$ was filled in $M$, every node $h_{j}^{1},h_{j}^{2}$ is filled by $M^{\prime}$.